OpenStudy (anonymous):

Which expression is equivalent to the given expression? (√48x^8y^4)^2/3 a.2^6√x b.4^6√x c. 4x^1/6 d.2x^3

OpenStudy (anonymous):

\[\huge\sqrt{48x^8y^4}^{2\over3}~?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh wait this is you again...my drug dealer, hows it goin buddy? :D

OpenStudy (anonymous):

Great , will you answer that question for a gram of ice?

OpenStudy (anonymous):

gram of ice..? -_-

OpenStudy (anonymous):

? meth

OpenStudy (anonymous):

\[\large\sqrt{48x^8y^4}^{2\over3}~?\]\[\large\sqrt[3]{48}^2~~\sqrt[3]{x^8}^2~~\sqrt[3]{y^4}^2\]\[\large(2\sqrt[3]{6})^2~~(2\sqrt[3]{x})^2~~(y\sqrt[3]{y})^2\] k im lost and this is confusing moreso because none of your answers have a \(y\)..

OpenStudy (anonymous):

i am just gonna guess lol, Thanks for you time

OpenStudy (anonymous):

yurp :D